OpenStudy (anonymous):

how would I solve 5^[x+2] + 5^x = 24

OpenStudy (anonymous):

Index Rule: \[5^{x+2} = 5^{x}\times5^{2}\]

OpenStudy (anonymous):

dont understand, what happened to the 24?

OpenStudy (dape):

Use the index rule to factor out the 5^x.

OpenStudy (anonymous):

\[(5^{2} \times 5^{x})+ 5^{x} = 24\]

OpenStudy (anonymous):

um okay, than what?

OpenStudy (dape):

So, when you factor out the 5^x you get: \[ 5^x(5^2+1)=5^x*26=24 \]

OpenStudy (anonymous):

\[5^{x} = \frac{ 24 }{ 26 }\]

OpenStudy (anonymous):

still confsued but thanks anyways

OpenStudy (dape):

Which step confuses you?

OpenStudy (anonymous):

Where you are confused?